leg1: x+2 leg2: x and the hypotenuse is square root of 10. how to find x using the pythagorean theorem?.
Answer by nerdybill(7384)

leg1: x+2 leg2: x and the hypotenuse is square root of 10. how to find x using the pythagorean theorem?.
.
Pythagorean theorem says:
(leg1)^2 + (leg2)^2 = (hypotenuse)^2
plug in what was given:
(x+2)^2 + x^2 = (sqrt(10))^2
(x+2)^2 + x^2 = 10
(x+2)(x+2) + x^2 = 10
x^2+4x+4 + x^2 = 10
2x^2+4x+4 = 10
2x^2+4x-6 = 0
x^2+2x-3 = 0
factoring:
(x+3)(x-1) = 0
x = {-3,1}
throw out the negative solution (extraneous) leaving:
x = 1